Strange Corporation purchased $60,400 of 6-year, 7% bonds of Kaecilius Inc. for $57,566 to yield an 8% return, and classified the purchase as an amortized cost method investment. The bonds pay interest semi-annually.

Required
Under ASPE and assuming the company has chosen the straight-line method of discount amortization - Prepare the following journal entries for Strange Corporation:

1) The purchase of the investment

2) Receipt of semi-annual interest and Discount amortization for the first interest payment that will be received

2) Receipt of semi-annual interest and Discount amortization for the second interest payment that will be received
Answer #1

--Working--

A Face Value $60,400
B Cash paid for Investment $57,566
C = A - B Discount on Bonds Investment $2,834
D Term [years] 6
E = C/D Annual Straight Line amortisation $472.33
F = E/2 Semi annual Amortisation $236.17

--Journal entries

Transaction Accounts title Debit Credit
1 Investment in Bonds $60,400.00
Discount on Bonds Investment $2,834.00
   Cash $57,566.00
(Purchase of investments)
2 Cash [60400 x 7% x 6/12] $2,114.00
Discount on Bonds Investment $236.17
Interest Revenue $2,350.17
(receipt of semi annual interest)
3 Cash [60400 x 7% x 6/12] $2,114.00
Discount on Bonds Investment $236.17
Interest Revenue $2,350.17
(receipt of semi annual interest)